# Learning to Explore in Diverse Reward Settings via Temporal-Difference-Error Maximization

### v1

@@ 1 Introduction
To this endeavor, we propose Stable Error-seeking Exploration (SEE), a novel approach for exploration that robustly handles diverse reward settings. SEE is a directed action selection mechanism that can be combined with any off-policy RL algorithm and does not introduce relevant additional hyperparameters. Our approach separates exploration from exploitation by optimizing for two disentangled objectives and training two individual policies. On the one hand, the exploitation objective remains the unchanged RL objective of maximizing the (discounted) cumulative return, and, thus, it can be combined with any existing RL optimization pipeline, including reward shaping mechanisms. On the other hand, our exploration objective maximizes the absolute temporal-difference-error (TD-error) encountered during training rollouts. The intuition behind this choice is that a high TD-error, especially in deterministic environments, indicates the presence of potentially relevant information not yet learned. We point out that decoupling exploration and exploitation into two separate objectives is not a new idea, as shown in prior works (Whitney et al., 2021; Simmons-Edler et al., 2020; Schäfer et al., 2022). However, till now, it has been challenging to devise effective and stable learning procedures when maximizing the TD-error. In this paper, we identify three distinct causes of instability affecting this setting, namely (i.) far off-policy learning, when a behavior policy is too out of distribution w.r.t. the target policy; (ii.) the conflict of interest of terminating an episode while the exploration objective follows an always positive reward signal; and (iii.) the non-stationary nature of the TD-error as a target. The contribution of this work is the formulation of a methodological solution to tackle each of these issues, i.e., (1.) combining the exploration and exploitation policies into a single behavior policy that bridges both distributions; (2.) using a maximum reward update, which is agnostic towards the length of an episode; (3.) conditioning the exploration value function on current estimates of the exploitation value function to inform it about the cause of change in the TD-error target. We integrate these solutions into one algorithmic approach, resulting in SEE.
@@ 3 Stable Error-seeking Exploration
We propose an off-policy exploration method that performs robustly in different reward scenarios without hyperparameter adjustments. To this end, we use separate objectives for exploitation and exploration. On the one hand, the exploitation objective remains the regular cumulative discounted return. On the other hand, the exploration objective aims to maximize the absolute value of the encountered TD-error of the exploitation objective throughout training rollouts. This choice entails using two separate policies, one for each objective. It is worth noting that prior works have already established this separation of objectives (Whitney et al., 2021; Simmons-Edler et al., 2020; Schäfer et al., 2022). However, despite its simple formulation, the exploration objective is challenging to optimize due to what we identify to be three main causes of instability in the learning process. In the following, we describe in detail these causes of instability and propose targeted solutions for each of them, eventually devising an algorithmic solution that mitigates them and enables stable and effective learning.
@@ 3.4 Exploration objective
The exploitation objective remains the unchanged objective of the underlying base RL algorithm, and it is only affected by the gathered data from the replay buffer during updates. Algorithm 1 shows how SEE can be combined with an arbitrary off-policy actor-critic RL method. For the sake of clarity, some details like target networks are omitted. We refer to Appendix D for implementation details.
@@ Acknowledgments
The authors also gratefully acknowledge the "Julia 2" HPC provided by the Julius-Maximilians-University Würzburg. "Julia 2" was funded as DFG project as "Forschungsgroßgerät nach Art 91b GG" under INST 93/1145-1 FUGG.
@@ Appendix C Hyperparameters
Hyperparameter | Classic | MuJoCo |
Learning rate | 0.001 | 0.0003 |
Batch size | 256 | 256 |
Discount | 0.99 | 0.99 |
Replay buffer size | ||
Target networks | ||
Update freq | ||
Warm-up steps | ||
Hidden dims | ||
Temperature | auto | auto |
Initial temperature | ||
@@
Hyperparameter | Classic | MuJoCo |
Learning rate | 0.001 | 0.001 |
Batch size | 256 | 256 |
Discount | 0.99 | 0.99 |
Replay buffer size | ||
Target networks | ||
Critic update freq | ||
Actor update freq | ||
Target update freq | ||
Warm-up steps | ||
Hidden dims | ||
Action noise std | ||
Target noise std | ||
@@
Hyperparameter | Classic | MuJoCo |
Action noise std | ||
@@
These were taken from the RL Baseline3 Zoo (Raffin, 2020). The parameter sets for the Classic environments were taken from the tuned version of Pendulum-v1 for TD3 and SAC respectively, and the MuJoCo sets were taken from HalfCheetah-v4 hyperparameters. SEE does not introduce relevant new hyperparameters but doubles them as we separately optimize for two objectives. All hyperparameters for the exploration objective are copied from the exploitation objective. For the FetchPickAndPlace environments, we use an additional hidden layer such that the hidden dims are . All remaining hyperparameters are the same as the MuJoCo setting.

### v2

@@ 1 Introduction
To this endeavor, we propose Stable Error-seeking Exploration (SEE), a novel approach for exploration that robustly handles diverse reward settings. SEE is a directed action selection mechanism that can be combined with any off-policy RL algorithm and does not introduce relevant additional hyperparameters. Our approach separates exploration from exploitation by optimizing for two disentangled objectives and training two individual policies. On the one hand, the exploitation objective remains the unchanged RL objective of maximizing the (discounted) cumulative reward, and, thus, it can be combined with any existing RL optimization pipeline, including reward shaping mechanisms. On the other hand, our exploration objective maximizes the absolute temporal-difference-error (TD-error) encountered during training rollouts. The intuition behind this choice is that a high TD-error, especially in deterministic environments, indicates the presence of potentially relevant information not yet learned. We point out that decoupling exploration and exploitation into two separate objectives is not a new idea, as shown in prior works (Whitney et al., 2021; Simmons-Edler et al., 2020; Schäfer et al., 2022). However, till now, it has been challenging to devise effective and stable learning procedures when maximizing the TD-error. In this paper, we identify three distinct causes of instability affecting this setting, namely (i.) far off-policy learning, when a behavior policy is too out of distribution w.r.t. the target policy; (ii.) the conflict of interest of terminating an episode while the exploration objective follows an always positive reward signal; and (iii.) the non-stationary nature of the TD-error as a target. The contribution of this work is the formulation of a methodological solution to tackle each of these issues, i.e., (1.) combining the exploration and exploitation policies into a single behavior policy that bridges both distributions; (2.) using a maximum reward update, which is agnostic towards the length of an episode; (3.) conditioning the exploration value function on current estimates of the exploitation value function to inform it about the cause of change in the TD-error target. We integrate these solutions into one algorithmic approach, resulting in SEE.
@@ 3 Stable Error-seeking Exploration
We propose an off-policy exploration method that performs robustly in different reward scenarios without hyperparameter adjustments. To this end, we use separate objectives for exploitation and exploration. On the one hand, the exploitation objective remains the regular cumulative discounted reward. On the other hand, the exploration objective aims to maximize the absolute value of the encountered TD-error of the exploitation objective throughout training rollouts. This choice entails using two separate policies, one for each objective. It is worth noting that prior works have already established this separation of objectives (Whitney et al., 2021; Simmons-Edler et al., 2020; Schäfer et al., 2022). However, despite its simple formulation, the exploration objective is challenging to optimize due to what we identify to be three main causes of instability in the learning process. In the following, we describe in detail these causes of instability and propose targeted solutions for each of them, eventually devising an algorithmic solution that mitigates them and enables stable and effective learning.
@@ 3.4 Exploration objective
The exploitation objective remains the unchanged objective of the underlying base RL algorithm. It is only affected by the gathered data from the replay buffer during updates. Algorithm 1 shows how SEE can be combined with an arbitrary off-policy actor-critic RL method. For the sake of clarity, some details like target networks are omitted. We refer to Appendix D for implementation details.
@@ Acknowledgments
The authors also gratefully acknowledge the "Julia 2" HPC provided by the Julius-Maximilians-University Würzburg. "Julia 2" was funded as DFG project as "ForschungsgroSSgerät nach Art 91b GG" under INST 93/1145-1 FUGG.
@@ Appendix C Hyperparameters
| Learning rate | 0.001 | 0.0003 | |||
| Batch size | 256 | 256 | |||
| Discount | 0.99 | 0.99 | |||
| Replay buffer size | |||||
| Target networks | |||||
| Update freq | |||||
| Warm-up steps | |||||
| Hidden dims | |||||
| Temperature | auto | auto | |||
| Initial temperature | |||||
@@
| Learning rate | 0.001 | 0.001 | |||
| Batch size | 256 | 256 | |||
| Discount | 0.99 | 0.99 | |||
| Replay buffer size | |||||
| Target networks | |||||
| Critic update freq | |||||
| Actor update freq | |||||
| Target update freq | |||||
| Warm-up steps | |||||
| Hidden dims | |||||
| Action noise std | |||||
| Target noise std | |||||
@@
| Action noise std | |||||
@@
These were taken from the RL Baseline3 Zoo (Raffin, 2020). The parameter sets for the Classic environments were taken from the tuned version of Pendulum-v1 for TD3 and SAC respectively, and the MuJoCo sets were taken from HalfCheetah-v4 hyperparameters. The discount factor is always set to . SEE does not introduce relevant new hyperparameters but doubles them as we separately optimize for two objectives. All hyperparameters for the exploration objective are copied from the exploitation objective. For the FetchPickAndPlace environments, we use an additional hidden layer such that the hidden dims are . All remaining hyperparameters are the same as the MuJoCo setting.